OpenStudy (anonymous):

(c) Riley also thought that the surface area would triple. Is Riley correct? How many times greater is the surface area of the larger aquarium than the smaller? Show your work and explain your reasoning.

OpenStudy (anonymous):

@PRAETORIAN.10 @demonchild99 @hysusonic @campbell_st @undeadknight26

OpenStudy (anonymous):

help me please

OpenStudy (anonymous):

i need helop

OpenStudy (anonymous):

@undeadknight26

undeadknight26 (undeadknight26):

Yes. Because i have no information to prove this...And today is yes day for me.

OpenStudy (anonymous):

@samanthagreer

OpenStudy (anonymous):

@TheRaggedyDoctor

OpenStudy (praetorian.10):

umm dont we need to know the specifics of what the surface area originally was so we know how to ensure it being trippled is correct (or am i just that sucky at math?)

OpenStudy (theraggedydoctor):

This looks like it's part of a different question? Can you post the rest?

OpenStudy (anonymous):

1. Riley noticed an aquarium in his father’s office. The aquarium had the dimensions 16 in. by 8.5 in. by 10.5 in. The formula for volume is: V = l × w × h.

OpenStudy (anonymous):

hate when they put ?'s

OpenStudy (praetorian.10):

ikr

OpenStudy (anonymous):

@undeadknight26 @PRAETORIAN.10 @demonchild99 @campbell_st @hysusonic

OpenStudy (hysusonic):

i was called?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

@HelpMePlease1234321

OpenStudy (anonymous):

@akonkel

OpenStudy (anonymous):

@Alieen

OpenStudy (anonymous):

i'm only taking 8th grade math sorry

OpenStudy (anonymous):

ok

OpenStudy (hysusonic):

what do you need help with now?

OpenStudy (anonymous):

2. Mr. Cooper’s company rents storage containers to customers. The customers can choose which size fits their needs and Cooper Storage delivers the containers. After the customers fill the containers, the company transports them to a warehouse to be stored until the customer needs the contents. (a) The volume for one size of the storage containers is 574 ft3. The container has a length of 10 ft and a width of 7 ft. How tall would a stack of 6 containers be if they were stacked one on top of another?

OpenStudy (anonymous):

blach wow it autocorrected to THAT?

OpenStudy (anonymous):

dudes idk the answer

OpenStudy (anonymous):

idk means i did know

OpenStudy (anonymous):

in the last question

OpenStudy (anonymous):

ok start with this: what is the formula to find volume of a rectangle?

OpenStudy (anonymous):

umm v = l x w x h

OpenStudy (anonymous):

So \[574 = 10 \times 7 \times h\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

wait H = 10.5?

OpenStudy (anonymous):

\[574= 70 \times h\]

OpenStudy (anonymous):

\[574 \div 70 = ?\]

OpenStudy (anonymous):

8.2?

OpenStudy (anonymous):

Yes! You got it!

OpenStudy (anonymous):

Any others?

OpenStudy (anonymous):

574 / 70 = 8.2 okaay

OpenStudy (anonymous):

so answer is 8.2?

OpenStudy (anonymous):

Yes the answer is 8.2

OpenStudy (anonymous):

(b) There are two other different sizes of containers. The price for renting a container per month for a residence within 30 miles is: 7’ container: 7 ft by 7 ft by 8 ft $99.00 12’ container: 12 ft by 8 ft by 8 ft $199.00 What is the price per cubic foot for each container? Round prices to the nearest cent. $6.00

OpenStudy (anonymous):

help me with this other one

OpenStudy (anonymous):

what is there for the question marks?

OpenStudy (anonymous):

container: 7 ft by 7 ft by 8 ft $99.00 12’ container: 12 ft by 8 ft by 8 ft $199.00 What is the price per cubic foot for each container? Round prices to the nearest cent. $6.00

OpenStudy (anonymous):

stupid qustion marks

OpenStudy (anonymous):

i know you have to delete the symbol and retype it

OpenStudy (anonymous):

@HelpMePlease1234321

OpenStudy (anonymous):

@undeadknight26

OpenStudy (anonymous):

i can do it i was typing =_=

undeadknight26 (undeadknight26):

smh Good job @Alieen <3

OpenStudy (anonymous):

\[7\times7\times8 = 392\]

OpenStudy (anonymous):

\[392 \div 99 =?\]

OpenStudy (anonymous):

0XUW8IJ9UESU,POVRSDHMVROIEMHTIHDVR

OpenStudy (anonymous):

Please don't spam, it cause lag and frequent crashes http://openstudy.com/code-of-conduct

OpenStudy (anonymous):

ok sorry

OpenStudy (anonymous):

3.95?

OpenStudy (anonymous):

@Alieen

OpenStudy (anonymous):

@undeadknight26

OpenStudy (anonymous):

@Nnesha

OpenStudy (anonymous):

help me please

OpenStudy (anonymous):

@CausticSyndicalist @Werewolfprincess

OpenStudy (anonymous):

@PRAETORIAN.10

OpenStudy (anonymous):

@DUES

OpenStudy (anonymous):

I'm sorry, Yes. It is 3.95 repeating

OpenStudy (anonymous):

@PreethaRam